# Computing dielectric spectra in molecular dynamics simulations: using a cavity to disentangle self and cross correlations

### v1

@@ I Introduction
where $\langle$.$\rangle$ . ⟩ is an ensemble average. Linking the molecular dynamics at the microscopic scale to the frequency-dependent dielectric permittivity $\epsilon(\omega)$, a macroscopic quantity, is not straightforward. This is due to the distinction between the applied electric field and the one felt by the microscopic volume. Taking this into account leads to [7, 8, 6]:
@@ II Methods
System Force Field $N$ $T$ (K) $\mu$ (D) $\lambda$
@@ III Static dielectric permittivity
where $V(r)/a^{3}$ is the volume fraction of the simulation box contained in the sphere of radius $r$. Knowing this contribution allows one to subtract it from $G_{\mathrm{K}}(r)$ and recover $g_{\mathrm{K}}(r)$. This is shown in Fig. 3 for water, propanol, and diethylether. Eq. 7 was evaluated without any adjustable parameter by computing $\epsilon$ from $G_{\mathrm{K}}$. We see that it reproduces perfectly the second increase of $G_{\mathrm{K}}(r)$ and, in particular, its maximum slope at $r=a/2$. For all three systems, $g_{\mathrm{K}}(r)$ quickly reaches its final value $g_{\mathrm{K}}$. This observation is consistent with the work on water of Zhang et al. [41, 56] who demonstrated that the effect of BCs could be removed by combining simulations equivalent to $\epsilon_{\mathrm{RF}}\rightarrow\infty$ and $\epsilon_{\mathrm{RF}}\rightarrow 0$.
Figure 4: Effect of the number of molecules $N$ in the simulation box for H2O and PrOH. The quantities shown are the same as in Fig. 3.
@@
The values of $g_{\mathrm{K}}$ and $\epsilon$ obtained this way are given in Table 2, and their convergence is shown in Figs. 2c and d in blue. We see that using a cavity to get $\epsilon$ gives a result perfectly compatible with the standard method relying on $G_{\mathrm{K}}$. It should be noted that the cavity method has some drawbacks: it requires simulating a large number of molecules so that the cavity remains small compared to the total system, which may be impractical for polarizable systems. Moreover, obtaining $G_{\mathrm{K}}(r_{\mathbf{c}})$ from MD trajectories requires much more computational power than $G_{\mathrm{K}}$. On the other hand, the cavity method appears to be much less sensitive to noise than the standard one. Indeed, the slow convergence of $G_{\mathrm{K}}$ seems to arise from $G_{\mathrm{BCs}}(r)$ rather than $g_{\mathrm{K}}(r)$. Consequently, accurate values of $\epsilon$ can be obtained from only several tens of total relaxation times instead of several hundred. This can be of interest when studying the dielectric response of liquids at lower temperatures, where relaxation times become much longer. Still, it has to be balanced with the larger simulation box required. For instance, in the present work, simulating a system containing 800 propanol molecules rather than 9600 is 5.5 times faster, which is not far from compensating for the gain in required simulation time. In practice, the benefits of the virtual cavity method become truly significant when a large simulation box is needed to avoid finite-size effects or if one wants to break down the cross-correlations as a function of distance to study how they build up. This is something we took advantage of in a previous work [32].
@@ IV Dielectric spectra
It should be noted that the radius $r$ is not as well defined as in the static case, as both molecules $i$ and $j$ have moved between time 00 and $t$. In practice we used $r_{ij}=|\vec{r}_{i}(0)-\vec{r}_{j}(0)|$ and we have checked for all three systems that evaluating the distance at time $t$ rather than at $t=0$ leads only to imperceptible changes in $C_{\mathrm{cross}}(r,t)$. The cross correlation functions corresponding to different times are shown for water as a function of $r$ in Fig. 5a. The already well visible time decorrelation is shown normalized for different radius $r$ in Fig. 5b, together with the self and the total box correlation functions. From each of these, we define a correlation time $\tau=\int_{0}^{\infty}C(t)/C(0)\mathrm{d}t$, shown for all three systems in Figs. 6a-c. We see that the dynamics of the cross correlations (green curve) is always slower than that of the self (blue dashed line) and depends on the radius $r$. It increases quickly at small $r$, reaching a maximum before decreasing slightly.
@@
The result is shown as dashed orange curves in Figs. 8d-f and appears reasonably close to the result of eq. 13 when focusing on the main relaxation peak for water and propanol. The main peak is the feature we are most interested in these non-polarizable simulations that are not able to reproduce accurately the high frequency part of dielectric spectra [49]. This is the method we used in previous work on glycerol [32, 46] and up to a prefactor it is equivalent to the approach chosen by Alvarez et al. [31] on water who consider a cavity of radius $r_{\mathrm{c}}=40$ in a giant simulation box ($a=333$). For diethylether, however, the approximation is not valid due to the dielectric permittivity being close to unity. This illustrates how the electrostatic BCs can affect the microscopic dynamics: the dipole moment of a small microscopic cavity decorrelates 10 % faster and less exponentially that the box total dipole moment which gives the correct dielectric spectra (see $C(t)/C(0)$ in Figs. 8a-c).
@@ VI Conclusion
We have studied in detail how dielectric spectra could be obtained in MD simulations with conductive electrostatic BCs by considering the fluctuation of the dipole in a virtual cavity rather than of the total simulation box. The cavity must be chosen large enough to contain the short-range dipolar correlation present in polar liquids, but small enough compared to the simulation box, so that correlations induced by the BCs remain negligible. We verified that this approach, in addition to being consistent with the standard method regarding the resulting dielectric spectra, provides direct access to the static and dynamic correlations representative of the experimental infinite case. In practice, it requires simulating a large system and performing a computationally costly analysis of the trajectories, but it has the benefit of requiring significantly less statistics to obtain dielectric spectra than the standard method. This is of practical interest when a very good signal-to-noise ratio is needed, for example, for accessing the non-linear dielectric response, which involves a subtraction between almost identical linear spectra [46].
The cavity method gives an occasion to shed a direct light on the effect of BCs in MD simulation, well known in the literature, but paradoxically concealed behind the simplicity of eq. 4. Due to the non-linear relationship between $\epsilon(\omega)$ and the molecular correlation function, there is no general, fully satisfactory way of decomposing a dielectric spectrum into self and cross parts that remains consistent with the infinite system case. Fortunately, this problem can be reasonably overcome when focusing on the main relaxation peak of polar liquids whenever $\Delta\epsilon>10$, which is of practical interest in such non-polarizable classical MD simulations. The cavity approach can then be used to study in details how short-range cross-correlations build-up to better understand their origin. Applying this method, in the case of 1-propanol, to investigate of how they are linked to supramolecular structures formed by hydrogen bonds [26] will be the subject of a future work.

### v2

@@ I Introduction
where $\langle.\rangle$ is an ensemble average. Linking the molecular dynamics at the microscopic scale to the frequency-dependent dielectric permittivity $\epsilon(\omega)$, a macroscopic quantity, is not straightforward. This is due to the distinction between the applied electric field and the one felt by the microscopic volume. Taking this into account leads to [7, 8, 6]:
@@ II Methods
System Force Field NN $T$ (K) $\mu$ (D) $\lambda$
@@ III Static dielectric permittivity
where $V(r)/a^{3}$ is the volume fraction of the intersection between a sphere of radius $r$ and the cubic simulation box. Knowing this contribution allows one to subtract it from $G_{\mathrm{K}}(r)$ and recover $g_{\mathrm{K}}(r)$. This is shown in Fig. 3 for water, propanol, and diethylether. Eq. 7 was evaluated without any adjustable parameter by computing $\epsilon$ from $G_{\mathrm{K}}$. It reproduces perfectly the second increase of $G_{\mathrm{K}}(r)$ and clearly illustrates that this effect is due to PBCs. Indeed, the flattening of $G_{\mathrm{K}}(r)$ for $r>a/2$ is related to $V(r)$ being the volume of a cube with side length $a$ truncated by a sphere of radius $r$ [38]. For all three systems, $g_{\mathrm{K}}(r)$ quickly reaches its final value $g_{\mathrm{K}}$. This observation is consistent with the work on water of Zhang et al. [41, 56] who demonstrated that the effect of BCs could be removed by combining simulations equivalent to $\epsilon_{\mathrm{RF}}\rightarrow\infty$ and $\epsilon_{\mathrm{RF}}\rightarrow 0$.
Figure 4: Effect of the number of molecules NN in the simulation box for H2O and PrOH on the $r$-dependent Kirkwood correlation factor directly computed from the simulation (a) and corrected from finite-size effects (b).
@@
The values of $g_{\mathrm{K}}$ and $\epsilon$ obtained this way are given in Table 2, and their convergence is shown in Figs. 2c and d in blue. We see that using a cavity to get $\epsilon$ gives a result perfectly compatible with the standard method relying on $G_{\mathrm{K}}$. It should be noted that the cavity method has some drawbacks: it requires simulating a large number of molecules so that the cavity remains small compared to the total system, which may be impractical for polarizable systems. Moreover, obtaining $G_{\mathrm{K}}(r_{\mathbf{c}})$ from MD trajectories requires much more computational power than $G_{\mathrm{K}}$. On the other hand, the cavity method appears to be much less sensitive to noise than the standard one. Indeed, the slow convergence of $G_{\mathrm{K}}$ seems to arise from $G_{\mathrm{BCs}}(r)$ rather than $g_{\mathrm{K}}(r)$. Consequently, accurate values of $\epsilon$ can be obtained from only several tens of total relaxation times instead of several hundred. This can be of interest when studying the dielectric response of liquids at lower temperatures, where relaxation times become much longer. Still, it has to be balanced with the larger simulation box required. For instance, in the present work, simulating a system containing 800 propanol molecules rather than 9600 is 5.5 times faster, which is not far from compensating for the gain in required simulation time. In practice, the benefits of the virtual cavity method become truly significant when a large simulation box is needed to avoid finite-size effects or if one wants to break down the cross-correlations as a function of distance to study how they build up. We previously took advantage of this [32] to reach low enough temperatures to separate the main relaxation peak from the THz range dynamics by up to six orders of magnitude, enabling precise characterization of the orientational dynamics.
@@ IV Dielectric spectra
It should be noted that the radius $r$ is not as well defined as in the static case, as both molecules $i$ and $j$ have moved between time 0 and $t$. One solution [53] is to consider, at each time $t$, all molecules $j$ within the sphere of radius $r$ around molecule $i$ ($r_{ij}=|\vec{r}_{i}(t)-\vec{r}_{j}(t)|$). In this case, molecules are allowed to enter or leave the virtual cavity, and their number can fluctuate. Another way is to consider only the molecules that were in the sphere at $t=0$ ($r_{ij}=|\vec{r}_{i}(0)-\vec{r}_{j}(0)|$), although some of them may be further apart than $r$ at time $t$. In both cases, the correlation function is then averaged over all molecules $i$ of the box and for initial times distributed over the total duration of the simulation. We have checked, for all three systems, that evaluating the distance at time $t$ rather than at $t=0$ leads only to negligible changes in $C_{\mathrm{cross}}(r,t)$ (see suppl. mat.), and we have chosen, in the following, to compute $r_{ij}$ at $t=0$. The cross correlation functions corresponding to different times are shown for water as a function of $r$ in Fig. 5a. The already well visible time decorrelation is shown normalized for different radius $r$ in Fig. 5b, together with the self and the total box correlation functions. From each of these, we define a correlation time $\tau=\int_{0}^{\infty}C(t)/C(0)\mathrm{d}t$, shown for all three systems in Figs. 6a-c. We see that the dynamics of the cross correlations (green curve) is always slower than that of the self (blue dashed line) and depends on the radius $r$. It increases quickly at small $r$, reaching a maximum before decreasing slightly.
@@
To reduce the computational cost of the simulation, we have neglected the effect of electronic polarizability, like many MD studies of this kind [22, 28, 31]. It is worth mentioning the consequences of this simplification. For water, polarizable and non-polarizable models have been directly compared in the literature [55, 45, 49], showing that taking into account polarizability is required to reproduce the experimentally observed features of dielectric spectra above $10^{12}$ Hz. In contrast, the main relaxation peak was shown to be not much affected by polarizability effects. When focusing on the structural relaxation, this makes non-polarizable models a reasonable trade-off by enabling larger systems to be simulated over longer periods.
@@
The result is shown as dashed orange curves in Figs. 8d-f and appears reasonably close to the result of eq. 13 when focusing on the main relaxation peak for water and propanol. This is the method we used in previous work on glycerol [32, 46] and up to a prefactor it is equivalent to the approach chosen by Alvarez et al. [31] on water who consider a cavity of radius $r_{\mathrm{c}}=40$ Å in a giant simulation box ($a=333$ Å). For diethylether, however, the approximation is not valid due to the dielectric permittivity being close to unity, and eq. 13 has to be used instead. This illustrates how the electrostatic BCs can affect the microscopic dynamics: the dipole moment of a small microscopic cavity decorrelates 10 % faster and less exponentially that the box total dipole moment which gives the correct dielectric spectra (see $C(t)/C(0)$ in Figs. 8a-c).
@@ VI Conclusion
We have studied in detail how dielectric spectra could be obtained in MD simulations with conductive electrostatic BCs by considering the fluctuation of the dipole in a virtual cavity rather than of the total simulation box. The cavity must be chosen large enough to contain the short-range dipolar correlation present in polar liquids, but small enough compared to the simulation box, so that correlations induced by the BCs remain negligible. We verified that this approach is consistent with the standard method regarding the resulting dielectric spectra for three liquids with dielectric permittivities ranging from low to high. In addition, it provides a direct access to the static and dynamic correlations representative of the experimental infinite case. In practice, it requires simulating a large system and performing a computationally costly analysis of the trajectories, but it has the benefit of requiring significantly less statistics to obtain dielectric spectra than the standard method. This is of practical interest when a very good signal-to-noise ratio is needed, for example, for accessing the non-linear dielectric response, which involves a subtraction between almost identical linear spectra [46].
The cavity method gives an occasion to shed a direct light on the effect of BCs in MD simulation, well known in the literature, but paradoxically concealed behind the simplicity of eq. 4. Due to the non-linear relationship between $\epsilon(\omega)$ and the molecular correlation function, there is no general, fully satisfactory way, when using either a virtual cavity or the total box dipole moment, of decomposing a dielectric spectrum into self and cross parts that remains consistent with the infinite system case. Fortunately, this problem can be reasonably overcome when focusing on the main relaxation peak of polar liquids whenever $\Delta\epsilon>10$, which is of practical interest in such non-polarizable classical MD simulations. The cavity approach can then be used to study in details how short-range cross-correlations build-up to better understand their origin. Applying this method, in the case of 1-propanol, to investigate of how they are linked to supramolecular structures formed by hydrogen bonds [26] will be the subject of a future work.
Supplementary Material
See the supplementary material (doi.org/10.60893/figshare.jcp.c.8004019.v1) for details on how the choice of method for defining a virtual cavity affects the computation of dynamical cross-correlations.